The average salary of merchandisers is $45,000 per year with a standard deviation of $6000.. What is the average salary below which 20% of the merchandisers earn? What is the average salary above which the top 10% of the merchandisers earn?
Answer by Boreal(15235)

z for 0.20 is -0.84
z=(x-mean)/sd
-0.84=(x-45000)/6000
-5040=x-45000
x=$39960
for top 10% z=1.28
1.28=(x-45000)/6000
7680=x-45000
x=$52680